# Field-free, Quasi-continuous Operation of Optical Nanofiber Interface with Two-dimensional Ferromagnetic Trap

### v3

@@ Appendix B A multi-scattering model for loading the 2D-MOT-ONF interface
The process of loading a magneto-optical trap is characterized by the time-dependent atom number and the characteristic size of the MOT, $N(t)$ and $\sigma(t)$ respectively. For a single-chamber MOT where atoms are directly captured from the background vapor, as in this work, the $N(t)$ dynamics follows quite accurately the differential equation $\dot{N}=R-\kappa N-\beta N^{2}/\sigma^{3}$. Here the loading rate $R$ depends on the laser and magnetic field parameters, as well as the background rubidium partial pressure $P_{\rm Rb}$. Instead, the single-body loss rate $\kappa$ is almost solely decided by the total pressure $P$. The two-body loss coefficient $\beta$ describes the light-assisted collisional losses [53]. For the setup in this work (Fig. 3) with $P\approx 10^{-6}~{}$ascal, the $\kappa\sim 10~{}{\rm sec}^{-1}$ at moderate $R$ limits the number of atoms in the MOT and correspondingly the two-body loss. We therefore expect
@@
We expect the Eqs. (1) (2) analysis to be valid for both 2D and 3D MOTs. Here, we assume the 2D-MOT size is isotropic in the $x-y$ plane, with $\sigma$ according to Eq. (2) and is $z-$independent. Furthermore, We assume the 2D-MOT length, $l=4~{}$mm, is time-independent. This is supported by the camera video (Fig. 3c) during the MOT loading. Physically, the stationary length is associated with the fact that the restoring force along $z$ in the (quasi)2D-MOT is too weak to support any interplay with the multi-scattering force.
@@
To facilitate the numerical fit for Fig. 3c, in Eq. (2) we phenomenologically set $\tilde{\sigma}(N,\rho_{0})={\rm max}(\sigma_{0},\sigma_{s})$, with $\sigma_{s}=\sqrt{N/(l\rho_{0})}$ for the density-limited 2D-MOT. The density limit $\rho_{0}$ is fixed quite accurately first, by comparing the ONF-coupled resonant optical depth for the same MOT sample (Fig. 4) with the theoretical estimation based on optical scattering in the near field [46, 37], leading to $\rho_{0}=4\times 10^{10}/{\rm cm}^{3}$. By this step, we assume that after the MOT is fully loaded, $\rho_{0}$ is approximately reached in the near field of ONF.
@@
We simply assume a uniform $g(r_{\perp})$ between $0\leq r_{\perp}\leq b$. With $b$ as auxiliary parameters, we fit the Fig. 3c data with Eq. (4) to retrieve $\kappa=(0.4~{}{\rm sec})^{-1}$. The fit suggests $\sigma_{0}\approx 0.6b$ for $t<0.1~{}$s, and then gradually expand to $\sigma(t)\approx b$ within one second. The ONF fluorescence is a local observable. Consequently, the atom number $N_{0}$ retrieved from the measurement depends on the $b$ parameter we choose in the model. For example, $N_{0}=3\times 10^{6}$ is reached by setting $b=100~{}{\bf\mu}$m.

### v5

@@ Appendix B A multi-scattering model for loading the 2D-MOT-ONF interface
The process of loading a magneto-optical trap is characterized by the time-dependent atom number and the characteristic size of the MOT, $N(t)$ and $\sigma(t)$ respectively. For a single-chamber MOT where atoms are directly captured from the background vapor, as in this work, the $N(t)$ dynamics follows quite accurately the differential equation $\dot{N}=R-\kappa N-\beta N^{2}/\sigma^{3}$. Here the loading rate $R$ depends on the laser and magnetic field parameters, as well as the background rubidium partial pressure $P_{\rm Rb}$. Instead, the single-body loss rate $\kappa$ is almost solely decided by the total pressure $P$. The two-body loss coefficient $\beta$ describes the light-assisted collisional losses [62]. For the setup in this work (Fig. 3) with $P\approx 10^{-6}$~Pascal, the $\kappa\sim 10~{\rm sec}^{-1}$ at moderate $R$ limits the number of atoms in the MOT and correspondingly the two-body loss. We therefore expect
@@
We expect the Eqs. (1) (2) analysis to be valid for both 2D and 3D MOTs. Here, we assume the 2D-MOT size is isotropic in the xyx-y plane, with $\sigma$ according to Eq. (2) and is $z$-independent. Furthermore, We assume the 2D-MOT length, l=4l=4~$m$, is time-independent. This is supported by the camera video (Fig. 3c) during the MOT loading. Physically, the stationary length is associated with the fact that the restoring force along $z$ in the (quasi)2D-MOT is too weak to support any interplay with the multi-scattering force.
@@
To facilitate the numerical fit for Fig. 3e, in Eq. (2) we phenomenologically set $\tilde{\sigma}(N,\rho_{0})={\rm max}(\sigma_{0},\sigma_{s})$, with $\sigma_{s}=\sqrt{N/(l\rho_{0})}$ for the density-limited 2D-MOT. The density limit $\rho_{0}$ is fixed quite accurately first, by comparing the ONF-coupled resonant optical depth for the same MOT sample (Fig. 4) with the theoretical estimation based on optical scattering in the near field [50, 39], leading to $\rho_{0}=3\times 10^{10}/{\rm cm}^{3}$. In this step, we use $\tilde{\Gamma}=2\pi\times 10$ MHz to account for the observed line broadening at the time (Appendix C). By this step, we assume that after the MOT is fully loaded, $\rho_{0}$ is approximately reached in the near field of ONF.
@@
We simply assume a uniform $g(r_{\perp})$ between $0\leq r_{\perp}\leq b$. With $b$ as auxiliary parameters, we fit the Fig. 3e data with Eq. (4) to retrieve $\kappa=(0.4~{\rm sec})^{-1}$. The fit suggests $\sigma_{0}\approx 0.6b$ for t<0.1t<0.1~s, and then gradueally expand to $\sigma(t)\approx b$ within one second. The ONF fluorescence is a local observable. Consequently, the atom number $N_{0}$ retrieved from the measurement depends on the $b$ parameter we choose in the model. For example, $N_{0}=2.5\times 10^{6}$ is reached by setting $b=100~{\bf\mu}$m.
Appendix C Details of linewidth measurements
C.1 Summary of linewidth measurements
As in Fig. 4a, we design a composite measurement sequence to probe the atomic gas through the 2D-MOT-ONF interface. Each measurement takes less than 600 ns (Fig. 4b), during which the cooling/repumping/heating beams are all shut off. These beams are then switched back on to ensure quasi-continuous operation of the 2D-MOT. To verify that cooling is effective between the nanosecond probes, our measurement results were repeated by increasing $T_{\rm rep}$ to 8 $\bf\mu$}s and 16 $\bf\mu$}s.
As outlined in Sec. III.5, we use two methods to retrieve the atomic response from the transient absorption $T(\Delta,t)$ probed at detuning $\Delta=\omega-\omega_{eg}$. Here $\omega$ and $\omega_{eg}$ are the laser frequency and the atomic transition frequency, respectively. The atomic transition for Fig. 4 data is $F=2-F^{\prime}=3$ hyperfine transition of 87Rb. For comparison, in this section, as in Fig. 7b, we also discuss the linewidth measurements for the nearby $F=2-F^{\prime}=2$ transition.
The first method to retrieve the transition linewidth follows the standard frequency scan [10, 17, 8] with repeated measurements. We estimate the steady-state optical depth at each detuning $\Delta$ as ${\rm OD}(\Delta)=-\ln\langle T(\Delta,t)\rangle_{t}$, using transmission data averaged over a time window at sufficiently late times $t\gg 1/\Gamma$. Specifically, for Fig. 4d, we average the data over the interval $t_{\rm start}<t<t_{\rm start}+5~{\rm ns}$, with $t_{\rm start}=195~{\rm ns}$, to generate the Fig. 4c data. A Lorentzian fit is then used to extract the line center and linewidth. We have verified during data analysis that the fitted lineshape is insensitive to variations in $t_{\rm start}$ for $t_{\rm start}>160~\mathrm{ns}$, within the $\delta\tilde{\Gamma}\sim 2\pi\times 100~\mathrm{kHz}$ fit uncertainty (Fig. 4c), suggesting that the steady-state response [47] is effectively reached for reliable retrieval of the linewidth. This assumption is also supported by single-atom-based simulations [50, 39]. Physically, after the weak probe is switched on, the non-adiabatic transients decay exponentially with $\tilde{\Gamma}t$. Accordingly, the systematic error associated with the transient response is exponentially suppressed, allowing accurate retrieval of the steady-state response [47].
The second method [63] is to fit the linewidth directly from a single-shot $T(\Delta,t)$ curve. This method is based on the observation that as long as the ONF-atom response is linear, then, the time-dependent transmission of the probe, such as the Fig. 4d data where $\Delta=0$, is decided by the linear spectral response via Fourier transform $\mathcal{F}$,
$E_{\rm out}(t)$ $=\mathcal{F}(E_{\rm in}(\Delta)h(\Delta)),$ (5)
$h(\Delta)$ $=e^{i\phi(\Delta)-{\rm OD}(\Delta)/2}.$
By modeling the complex phase shift $\varphi(\Delta)=\phi(\Delta)+i{\rm OD}(\Delta)/2$ [64] via Beer-Lambert law [45], $\varphi\propto\rho\alpha l$ and $\alpha\propto 1/(\Delta-i\Gamma/2)$, it is possible to retrieve the atomic information directly by fitting the single-shot transmission data via $T(t)=|E_{\rm out}(t)/E_{\rm in}(t)|^{2}$. As detailed in ref. [63], to efficiently retrieve the atomic response, the probe envelope $E_{\rm in}(t)$ must be complex enough. Therefore, instead of using a simple square pulse (Fig. 4a), we keep the probe on for t>200t>200~ns, but with its frequency shifted by 250 MHz (not shown in Fig. 4). An interference is induced between the coherent forward emission – see Fig. 4b at t>200t>200~ns for the intensity profile of the “flash” [40, 41] – with the frequency-shifted probe. The interference data is complex enough for us to infer the transition center and linewidth within a single shot [63].
We verify that the frequency-scan method and the single-shot method agree. The single-shot method, which utilizes the full $T(t)$ profile, is more data-efficient and is exploited to investigate the origin of the line broadening observed in Fig. 4(c,d). In the following, we present evidence that the line broadening is related to the surface quality of the ONF [49] and is likely of magnetic origin.
C.2 On magnetic broadening of absorption profile
The absorption line of hyperfine transition can be efficiently broadened by a magnetic field, if it is perpendicular to the quantization axis set by light. For example, for the $F=2-F^{\prime}=3$ transition of 87Rb probed by an $x$-polarized light, the transition frequency of the $\sigma^{\pm}$ transitions along $z$ is split by as much as 1.9 MHz per Gauss [47]. This splitting of transition frequency efficiently broaden the overall absorption profile. Assuming equal population on the $F=2$ Zeeman sublevels, numerical simulation [50, 39] for the $F=2-F^{\prime}=3$ transition suggests a 1-Gauss field lead to $\sim$1 MHz line broadening while largely keeping the Lorenzian shape. Larger field does distort the line, but we expect that when the field strength has a distribution, then the distortion can be somewhat smoothed out. The broadening is less efficient for the cases where the light-atom interaction commute with Zeeman shifts, such as the $\pi$ or $\sigma$ transition alone.
The line-broadening effect by the longitudinal field, as described above, depends on the magnetic $g_{F}$ factors as well as the distribution of $\sigma^{\pm}$ transition strengths between Zeeman sublevel [47]. Our numerical simulation suggests that the broadening is approximately halved for the $F=2-F^{\prime}=2$ transition. This difference of magnetic response motivates us to investigate the two transitions together, as to be detailed in the following.
C.3 MOT alignment
Optimal operation of the 2D-MOT requires balancing the intensities of counter-propagating MOT beams [20]. Here, for the prism-MOT as in Fig. 1c and Fig. 3d, we achieve the balance by shifting the position of main beam and walk the angles of the small beams. The balance is verified by the insensitivity of the MOT center, according to the direct and prism-reflected ONF-2D-MOT images (Fig. 3(b,c)), following a 5-fold reduction of the MOT beam power. The slimmer 2D-MOT at the lower MOT beam power ($\sim$1 mW/cm2 per beam at the $\Delta/2\pi=-12$ MHz MOT detuning) also helps us to verify that the ONF is aligned along the zero-field line to within 50 mrad, at least.
C.4 Magnetic field scan
Figure 7: (a) The measured linewidth of $F=2-F^{\prime}=3$ transition as a function of the $B_{z}$ offset. (b) The measured linewidth of $F=2-F^{\prime}=2$ transition as a function of the $B_{z}$ offset. (c) The evolution of the $F=2-F^{\prime}=3$ linewidth following an increase of the heating laser power (which is shut off during probe) in a few hours.
We use the plate-coil and $x$-coil as in Fig. 3d to shift the $B$-field bias in the xyx-y plane. We scan the field by up to 1 Gauss in each direction, hardly see a change of linewidth, before the observed absorption signal is too small to support reliable readouts. This is expected. As already suggested by the Fig. 3f data, the shift of the MOT center rapidly reduces the number of atoms interacting with ONF in the near field. We therefore keep the $B_{x,y}$ offsets that optimize the optical depth experienced by the ONF-coupled resonant probe, in all the other measurements.
We use an additional $z$-coil with diameter $D_{z}=200$~$m$, mounted $\sim$300 mm away, to generate a $B_{z}$-offset. Similar to the case of $x$-coil, we use Radia [29, 30] to confirm that when the $B_{z}$ offset is generated, the field variation along the $l=4$ mm ONF interaction length is below $1\%$. With the $B_{z}$ offset, we repeat the single-shot linewidth measurement as by Eq. (5[63]. Typical results on the retrieved linewidth, for the case of linearly polarized HE11 probe (Appendix C.7), are presented in Fig. 7a. For circularly polarized HE11 probe, we also find shift of line center at MHz per Gauss level, as expected. Near $B_{z}=0$, however, the minimal linewidth $\tilde{\Gamma}/2\pi$ is as large as 10.5 MHz. This linewidth is the smallest we can get at the time of the measurement (See Appendix C.5). We verified that this minimal linewidth is not sensitive to the polarization state of the HE11 probe.
C.5 Drift of linewidth and heating-induced recovery
[…]
C.6 $F=2-F^{\prime}=2$ transition
As in Fig. 7b, we also perform the $F=2-F^{\prime}=2$ linewidth measurements during the $B_{z}$ scan. Comparing to the $\Gamma=2\pi\times 6.1$~MHz linewidth, the amount of broadening for the $F=2-F^{\prime}=2$ line is roughly half of that for the $F=2-F^{\prime}=3$ line in Fig. 7a. This agrees with the Appendix C.2 analysis, suggesting that the line broadening is of magnetic origin.
C.7 Probe intensity and polarization
The optical power of the probe pulse is calibrated by a single-photon detector (Excelitas SPCM-AQRH-16) to be $P_{\rm probe}=5$~pW [44]. By doubling the probe power and repeating the measurements for consistency, we verified that the probe is well below saturation in the near field.
The probe polarization is controlled by a $1/4$ plate and a $1/2$ plate right before the fiber coupling [65]. The polarization status of the HE11 mode at the ONF section is measured by observing Rayleigh scattering [66] through direct and mirror-reflected images in our prism-MOT setup (Fig. 3b). By scanning the motorized-stage-mounted waveplates and read the Rayleigh scattering along both $x,y$ directions, we are able to model the fiber polarization status quite accurately. This facilitates probe measurements with different HE11 polarizations, as in Appendix C.4.
C.8 Other factors
Our probe laser (Moglabs CEL) has a linewidth below 1 MHz. The reliable operation of the OAWG laser probe [67] is verified with independent measurements [64]. The atomic density in the ONF near field is expected to be moderate. By varying the MOT cooling power to reduce the atomic sample size, we verified that the linewidth is consistent for small (OD=0.6) and large (OD=2) optical depth.
C.9 Discussions
In this Appendix, we have detailed linewidth measurement in our ferromagnetic 2D-MOT-ONF system. Exploiting the field-free interface, the measurement at $f_{\rm rep}=250$~kHz repetition rate helps us to obtain good signal statistics within minutes. By the fast probe calibration, as presented in Fig. 4(a,b), our ONF-spectroscopy is robust against low-frequency technical noises. Exploiting the nanosecond transient probe (also see ref. [63]), our measurements preserve the atomic state and density distribution in the near field. That we see a broadened atomic line without substantial surface-interaction induced asymmetry [15, 49, 17, 8] is quite unexpected.
The data presented in Fig. 7 provide key evidence that the line-broadening is not only of magnetic origin but also affected by the heating laser through ONF. Up to the completion of this work, we do not have a definite conclusion on the origin of the line broadening. Nevertheless, from Fig. 7 we speculate that the broadening could be induced by certain nanoparticles that are magnetic, which are assembled to the ONF surface during prolonged ONF operation in moderate vacuum. Previous study have suggested that atoms adsorbed onto the nanofiber surface may assemble into nano-structures to affect the surface interaction [49].